OpenStudy (anonymous):

which equation shows the substitution method being used to solve the system of linear equations x+y=6 and x=y+5

OpenStudy (anonymous):

x + y = 6 x = y + 5 --- by substitution: x + y = 6 (y + 5) + y = 6 --- answer: C. (y + 5) + y = 6 ---

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

@michelleperez